5 (x-2)+6=2x+3 (x+1)
Answer by stanbon(75887)

5 (x-2)+6=2x+3 (x+1)
---------
5x-10 = 2x + 3x + 3
------
5x-10 = 5x + 3
----
That is a contradiction::
No solution
---------
Cheers,
Stan H.